A new inventory management system for the ABC Company could be developed at a cost of $280,000. The estimated net operating costs and estimated net benefits over six years of operation would be:

Year Estimated Net Estimated Net Benefits
Operating Costs
0 $280,000 $0
1 8,000 42,000
2 9,500 78,000
3 12,000 82,000
4 13,000 115,000
5 15,500 120,000
6 25,500 140,000

a. What would the payback period be for this investment? Would it be a good or bad investment? Why?

b. What is the ROI for this investment?

c. Assuming a 16 percent discount rate, what is this investment’s NPV?
Answer #1

The net cash inflows are calculated as:

2 3 4 Year Operating Costs -280000 -8000 -9500 Net 0 42000 78000 - 12000 82000 70000 - 13000 115000 102000 -15500 120000 1045

a. Payback period is the time taken to recover the initial investment.

Here cumulative cash inflow till 4th year = 274,500, So the company has to earn 5,500 more to recover initial investment of 280,000.

Cumulative cash inflow till 5th year = 379,000

Therefore the payback period lies between 4th & 5th year.

Difference in cumulative cashflows between 4th & 5th year = 104,500 (in 1 year)

Time taken to earn 5500 = 5,500/104,500= 0.053 years

Therefore Payback period = 4+0.053 = 4.053 years.

The earlier the payback period, better the investment. Here the payback period is after 4 years which is a long time. So this is a not so good investment. Also, we have to consider other parameters like NPV & IRR also. If the Internal rate of return is greater than required rate of return, the investment is considered to be good.

b. ROI is the same as Internal rate of return of the investment

IRR is calculated as follows:

C D E F G H 1 14 Year Operating 15 Costs 0 -280000 -8000 2 -9500 78000 68500 3 | -12000 82000 70000 4 -13000 115000 102000 5

IRR = 16%

c. Discount rate given to calculate NPV is 16% which is equal to the IRR.

We know that at IRR, NPV = 0. Hence the answer